Solve the quadratic equation using the square root property.
(x+8)^2 = -49

Answer by mananth(16949)

(x+8)^2= -49
(x+8)^2= -1*7^2
take square root
x+8 = +/- sqrt(-1)*7
x=-8+/- 7i